6. Bob deposited ₱25,000 in a new savings account at 6% interest compounded semiannually.


At the beginning of year 3, Bob deposits an additional ₱35,000 at 8% interest compounded semiannually. At the end of 4 years, what is the balance in Bob's account?




1
Expert's answer
2022-01-13T17:31:31-0500

Compound interest


"P=P_0(1+\dfrac{r}{n})^{nt}"

Bob deposited ₱25,000 in a new savings account at 6% interest compounded semiannually.

At the end of 4 years


"P_1=25000(1+\dfrac{0.06}{2})^{2(4)}=31669.2520"

At the beginning of year 3, Bob deposits an additional ₱35,000 at 8% interest compounded semiannually.

At the end of 4 years


"P_2=35000(1+\dfrac{0.08}{2})^{2(2)}=40945.0496"


"P=P_1+P_2=31669.2520+40945.0496"

"=72614.30"

At the end of 4 years, the balance in Bob's account is "\u20b172,614.30."